PROBING VIBRATIONALLY EXCITED STATES OF CH_{3}OOH THROUGH ACTION SPECTROSCOPY

\underline{JAMIE MATTHEWS} and AMITABHA SINHA, Department of Chemistry and Biochemistry, University of California-San Diego, 9500 Gilman Drive, La Jolla CA 92093-0314.

State selected photodissociation is used to probe the OH-stretching overtone states and corresponding torsional combination bands of CH_{3}OOH at room temperature and in a jet. The action spectra are analyzed with the aid of ab initio calculations and provide information about changes in rotational constant, torsional barrier height and equilibrium torsional angle with OH-stretching excitation. We also find that the OH fragment vibrational state distribution depends on the parent CH_{3}OOH vibrational level that is selected, with the OH stretch-HOO bend combination band producing substantially less OH(v=1) fragments compared to OH-stretching overtone states. These differences in OH(v=1) yield likely reflect enhanced vibrational state mixing associated with HOO bending mode excitation in CH_{3}OOH compared to excitation of zeroth order pure OH stretching states.